OpenStudy (anonymous):

Find the equation of the line that is perpendicular to the line formed by the equation 3x + 2y = -14 and passes through the point (-48, 3).

OpenStudy (anonymous):

first change the equation to the standard form of y = mx + b where m is the slope of the line 2y = -3x - 14 3 y = - ----x - 7 2 now it is in the standard form ans so we get slope i.e. m = -3/2 if m2 is the slope of the perpendicular line, then 1 1 m2 = - ------ = - ----- = 2/3 m -3/2 so the perpendicular line will be of the form y = m2x + b 2 => y = ---x + b ---------------(1) 3 we r given the point on the perpendicular line as (-48,3) i.e. x = -48 and y = 3 substituting the values in (1) we get 2 => 3 = ---- (-48) + b 3 Multiplying by 3 throughout => 9 = 2(-48) + 3b => 9 = -96 + 3b => 9 + 96 = 3b => 105/3 = b => 35 = b Putting this in (1), equation of perpendicular line is 2 y = ---x + 35 3